A seedman has seeds worth $0.50 per pound mixed with seeds worth $0.80 per pound. If he wishes to make a mixture of 30 lbs to sell at $0.75 per pound, how many pounds of each should he use?
5 lbs of $0.50
25 lbs of $0.80

Answer by checkley71(8403)

.8X+.5(30-X)=.75*30
.8X+15-.5X=22.5
.3X=22.5-15
.3X=7.5
X=7.5/.3
X=25 POUNDS OF $.80 MIX IS USED.
30-25=5 POUNDS OF $.50 MIX IS USED.
PROOF.
.8*25+.5*5=22.5
20+2.5=22.5
22.5=22.5